OpenStudy (anonymous):

find the x intercept of f(x)= 2x^3-15x^2+24x+19?

OpenStudy (anonymous):

\[1/2 (5-(3 3^(2/3))/(11-2 \sqrt(10))^(1/3)-(3 (11-2 \sqrt(10)))^(1/3)) \]

OpenStudy (anonymous):

x~~-0.571767

OpenStudy (anonymous):

x = 1/2 (5-(3 3^(2/3))/(11-2 sqrt(10))^(1/3)-(3 (11-2 sqrt(10)))^(1/3))

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

just tried, but i still dont get it.. what is the formula you used?

OpenStudy (anonymous):

This link will take you directly to the root finding portion, no need to scroll any further

OpenStudy (anonymous):

alright, thanks a lot!

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

By the way, whats your present level?

OpenStudy (anonymous):

I mean is horner scheme in the scope of your syllabus

OpenStudy (anonymous):

No its not.. but i ll try find some other websites to guide me through.. thanks for the help anyway!